OpenStudy (anonymous):

Check my answer? Simplify. (2^2)^-3 -64

OpenStudy (campbell_st):

nope its\[2^{2 \times -3} = 2^{-6} = 1/2^6 = 1/64\]

OpenStudy (anonymous):

Oh ok I know what I did wrong. Thanks